# Diffusion Feedback Helps CLIP See Better

### v3

@@ 3.2 Overall Structure of DIVA
As illustrated in Fig. 2, DIVA mainly consists of two parts: the CLIP model to be enhanced in terms of visual perception capabilities, and the pre-trained diffusion model providing generative feedback. Taking an original image and an empty text (as indicated by "Null" in Fig. 2) as inputs, the CLIP model encodes the corresponding visual features, which will be added with the empty text’s embeddings from the diffusion model’s text encoder for diffusion’s condition. Given the image with added noise, the diffusion model attempts to predict the noise added from the previous step to the current step with the aforementioned condition. This process needs to be repeated N times because, for each image, we will randomly select N states (i.e., two) from the total steps (e.g., 0-1000 steps) of the diffusion model for optimization. The corresponding loss function can be represented as Equation 5. Keeping all parts’ weights except the CLIP model frozen, the training objective is simply to minimize the reconstruction loss (i.e., generative guidance). In this manner, by constraining the diffusion model to more accurately predict the added noise, the CLIP’s original semantic-rich discriminative representations will be gradually optimized into representations with more visual details through diffusion feedback. On this basis, results in Sec. 4.4 demonstrate that our DIVA does not significantly damage the zero-shot performance of the original CLIP models. The pseudo code of the specific enhancement process can be found at Algorithm 1 in Appendix.
@@ 3.3 Diffusion’s Condition Design
Visual Recap Density. Although the visual dense recap scheme appears straightforward, the density of the recap is crucial. If the density is too high (i.e., introducing too many local tokens), the richness of the condition information approaches its maximum, greatly reducing the difficulty of reconstruction task. This results in CLIP’s representation requiring minimal optimization to easily complete the reconstruction, limiting the upper bound of CLIP’s optimized capabilities. Conversely, if the recap density is too low (i.e., retaining only class token or introducing few local tokens), CLIP’s optimization process will struggle with the high difficulty of reconstruction, failing to adequately learn the expected detailed visual representations. This intuitive point is confirmed in Sec. 4.5. Specifically, ensuring the visual class token is always present in the condition, we introduce randomly selected local token features with approximately 15% and 30% probabilities for OpenAI CLIP (Radford et al., 2021b) at 224 and 336 resolutions. For the SigLIP ViT-SO-14 (Zhai et al., 2023) at 224 and 384 image sizes, we incorporate local token features obtained through average pooling with local window sizes of 1$\times$6 and 1$\times$10, respectively. Except that introducing 50% randomly selected patch tokens into the condition for DFN ViT-H-14/378, for the remaining baselines (Fang et al., 2023; Xu et al., 2023a), we include all local token features as part of the condition. Apart from DFN ViT-H-14/224 and SigLIP ViT-SO-14/224&384 (Zhai et al., 2023) only using visual class token, all other models incorporate local features consistent with the training stage conditions during inference, combining them with class token to fully leverage the detailed representations captured by the enhanced CLIP.
@@ 4.1 Implementation Details
DIVA is trained on 8 NVIDIA-A100 80GB GPUs with a global batch size of 640. We adopt Stochastic Gradient Descent (SGD) optimizer with a learning rate of 1e-4 and momentum of 0.9 to refine the representations of CLIP models via generative feedback. We only optimize the CLIP models with relatively high-quality Conceptual-3M dataset (Sharma et al., 2018) for 4600 steps (i.e., nearly 1 epoch) among the training phase, which can already boost CLIP’s performance in a training-efficient manner. For all experiments, we adjust all the parameters of the discriminative CLIP models and keep the diffusion models’ parameters frozen through the whole training process.
@@ 4.5 Ablation Study
We conduct ablation studies on MMVP-VLM regarding each pattern and average score, exploring the effect of condition designing, introduced data scale and adopted diffusion models for DIVA.

### v4

@@ 3.2 Overall Structure of DIVA
As illustrated in Fig. 2, DIVA mainly consists of two parts: the CLIP model to be enhanced in terms of visual perception capabilities, and the pre-trained text-to-image diffusion model providing generative feedback. Taking an original image as input, the CLIP model encodes the corresponding visual features, which will be combined with the empty text’s embeddings (i.e., [BOS] & [EOS]) from the diffusion model’s text encoder for diffusion’s condition. Given the image with added noise, the diffusion model attempts to predict the noise added from the previous step to the current step with the aforementioned condition. This process needs to be repeated N times because, for each image, we will randomly select N states (i.e., two) from the total steps (e.g., 0$\sim$1000 steps) of the diffusion model for optimization. The corresponding loss function can be represented as Equation 5. Keeping all parts’ weights except the CLIP visual encoder frozen, the training objective is simply to minimize the reconstruction loss (i.e., generative guidance). In this manner, by constraining the diffusion model to more accurately predict the added noise, the CLIP’s original semantic-rich discriminative representations will be gradually optimized into representations with more visual details through diffusion feedback. On this basis, results in Sec. 4.4 demonstrate that our DIVA does not greatly damage the zero-shot performance of the original CLIP models. The pseudo code of the specific enhancement process can be found at Algorithm 1 in Appendix.
@@ 3.3 Diffusion’s Condition Design
Visual Recap Density. Although the visual dense recap scheme appears straightforward, the density of the recap is crucial. If the density is too high (i.e., introducing too many local tokens), the richness of the condition information approaches its maximum, greatly reducing the difficulty of reconstruction task. This results in CLIP’s representation requiring minimal optimization to easily complete the reconstruction, limiting the upper bound of CLIP’s optimized capabilities. Conversely, if the recap density is too low (i.e., retaining only class token or introducing few local tokens), CLIP’s optimization process will struggle with the high difficulty of reconstruction, failing to adequately learn the expected detailed visual representations. This intuitive point is confirmed in Sec. 4.5. Specifically, ensuring the visual class token is always present in the condition, we introduce randomly selected local token features with approximately 15% and 30% probabilities for OpenAI CLIP (Radford et al., 2021b) at 224 and 336 resolutions. For the SigLIP ViT-SO-14 (Zhai et al., 2023) at 224 and 384 image sizes, we incorporate local token features obtained through 1D average pooling with local window sizes of 6 and 10, respectively. Except that introducing 50% randomly selected patch tokens into the condition for DFN ViT-H-14/378, for the remaining baselines (Fang et al., 2023; Xu et al., 2023a), we include all local token features for the condition design. Apart from DFN ViT-H-14/224 and SigLIP ViT-SO-14/224&384 (Zhai et al., 2023) only using visual class token, all other models incorporate local features consistent with the training stage conditions during inference, combining them with class token to fully leverage the detailed representations captured by the enhanced CLIP.
@@ 4.1 Implementation Details
DIVA is trained on 8 NVIDIA-A100 80GB GPUs with a global batch size of 640. We adopt Stochastic Gradient Descent (SGD) optimizer with a learning rate of 1e-4 and momentum of 0.9 to refine the representations of CLIP models via generative feedback. We only optimize the CLIP models with relatively high-quality Conceptual-3M dataset (Sharma et al., 2018) for 4600 steps (i.e., nearly 1 epoch) among the training phase, which can already boost CLIP’s performance in a training-efficient manner. For all experiments, we adjust the parameters of the discriminative CLIP vision encoders and keep the pre-trained diffusion models’ parameters frozen through the training process.
@@ 4.5 Ablation Study
By taking OpenAI ViT-L-14/224 as the baseline model, we conduct comprehensive ablation studies on MMVP-VLM regarding each pattern and average score, exploring the effect of condition designing, introduced data scale and adopted diffusion models for DIVA.